The moment of inertia of a meter scale of mass 0.6kg about an axis perpendicular to the scale and located at the 20cm position on the scale in kgm² is?

Grade:11

1 Answers

Arun
3 years ago
Using the Parallel Axis Theorem,
 
the Moment of Inertia about an axis at an distance of 20 cm from its centre is given as,
 
  I = ML²/12 + Ma²
 
Now,
 
Using the conditions, given in Questions,
 
  M = 0.6 kg, L = 1 m.
 
a = 20/100 m = 0.2 m.
 
Substituting these values in the Formula, we will get,
 
I = 0.6 (1²/12 + 0.2²)
 
∴ I = 0.6(1/12 + 4/100)
 
∴ I = 0.6(0.0833 + 0.04)
 
∴ I = 0.6 × 0.123
 
∴ I = 0.074 kg-m².